OpenStudy (anonymous):

giving medal 3(2y - 0.3) = 19.4 - y a. 4.7 c. 3.8 b. 2.9 d. 10.1

OpenStudy (anonymous):

B :)

OpenStudy (anonymous):

@thmdkhs it's B. :)

OpenStudy (anonymous):

thnx!!

OpenStudy (anonymous):

No problem. :) Glad I could help. @thmdkhs

OpenStudy (anonymous):

3(2y - 0.3) = 19.4 - y open the bracket by multiplying the 3 with everything inside it: 6y - 0.9= 19.4-y move the -y , by adding +y to both sides: 6y - 0.9 +y= 19.4 move the -0.9 by adding +0.9 to both sides: 6y+y=19.4+0.9 add everything: 7y=20.3 remove the 7 by dividing by 7: 7y/7=20.3/7 you'll get: y=20.3/7= 2.9 The answer is: y=2.9